Circular No. 8921

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(153591) 2001 SN_{263}

M. C. Nolan and E. S. Howell, Arecibo Observatory; L. A. M. Benner, S. J. Ostro, and J. D. Giorgini, Jet Propulsion Laboratory; M. W. Busch, California Institute of Technology; L. M. Carter and R. F. Anderson, Smithsonian Institution; C. Magri, University of Maine at Farmington; D. B. Campbell and J. L. Margot, Cornell University; R. J. Vervack, Jr., Applied Physics Laboratory, Johns Hopkins University; and M. K. Shepard, Bloomsburg University, report that Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm), obtained on 2008 Feb. 12 and 13, show that minor planet (153591) is a triple system. Based on range extents at 75-m resolution, preliminary estimates of average diameters are 2 km, 1 km, and 400 m for the three components. The orbital separation for the larger two components is at least 10 km.

COMETS C/2007 Y3–Y9 (SOHO) Further to IAUC 8920, additional near-sun comets have been found

on SOHO website images; all were Kreutz sungrazers except for Marsdengroup member C/2007 Y4 and Meyer-group member C/2007 Y8 — both of which were small and stellar in appearance, reaching mag ∼ 6.5 (though the latter object brightened extremely fast in the first hour or two of visibility and became noticeably larger in that time — atypical for Meyer-group objects). K. Battams suggested the identity of C/2007 Y4 with C/2002 R4 (cf. MPEC 2002-S35; IAUC 7984), and B. G. Marsden published a linked orbit on MPEC 2008-B49 that suggests an approach to within 0.063 AU of the earth on 1997 June 14.7 (T = 1997 May 15.2). C/2007 Y3 was quite diffuse and small (mag ∼ 7.5). C/2007 Y5 was quite bright (mag ∼ 5.5) and teardrop-shaped (with a hint of a stubby tail). C/2007 Y6 was very diffuse (mag ∼ 8) with a hint of a tail. C/2007 Y7 was tiny and a little diffuse (mag ∼ 8.5). C/2007 Y9 was quite diffuse, only reaching mag 7.5.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 Y3 Dec. 21.104 17^{h}58.^{m}9 −25^{o}14′ C2 SY 2008-B49
C/2007 Y4  21.754 18 03.4 −24 07 C2 RK 2008-B49
C/2007 Y5  22.571 18 13.9 −28 25 C3 AK 2008-B49
C/2007 Y6  26.338 18 23.2 −25 00 C2 SY 2008-B49
C/2007 Y7  26.688 18 25.0 −24 47 C2 HS 2008-B49
C/2007 Y8  28.079 18 25.6 −22 10 C2 RK 2008-B61
C/2007 Y9  28.638 18 37.4 −26 06 C3 RM 2008-B61
2008 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green